# Digital-physical testbed for ship autonomy studies in the Marine Cybernetics Laboratory basin

### v1

@@ Abstract
The algorithms developed for Maritime Autonomous Surface Ships (MASS) are often challenging to test on actual vessels due to high operational costs and safety considerations. Simulations offer a cost-effective alternative and eliminate risks, but they may not accurately represent real-world dynamics for the given tasks. Utilizing small-scale model ships and robotic vessels in conjunction with a laboratory basin provides an accessible testing environment for the early stages of validation processes. However, designing and developing a model vessel for a single test can be costly and cumbersome, and often researchers lack availability to such infrastructure. To address these challenges and enable streamlined testing, we have developed an in-house testbed that facilitates the development, testing, verification, and validation of MASS algorithms in a digital-physical laboratory. This infrastructure includes a set of small-scale model vessels, a simulation environment for each vessel, a comprehensive testbed environment, and a digital twin in Unity. With this, we aim to establish a full design and verification pipeline that starts with high-fidelity simulation models of each model vessel, to the model-scale testing in the laboratory basin, allowing possibilities for moving to semi-fullscale validation with the R/V milliAmpere 1 passenger ferry and full-scale validation using the R/V Gunnerus. In this work, we present our progress on the development of this testbed environment and its components, demonstrating its effectiveness in enabling ship guidance, navigation, and control (GNC) including autonomy.
@@ 1 Introduction
Autonomous vessels have strong potential to significantly influence the maritime sector by enhancing safety, operational efficiency, and economic viability (gu2021autonomous; munim2019autonomous). Recent advances in sensor technology, artificial intelligence, and marine cybernetics have spurred the development of prototype autonomous vessels of various sizes (schiaretti2017survey). Additionally, large-scale initiatives have emerged to demonstrate the real-world feasibility of uncrewed ships under diverse conditions (reddy2019zero). These efforts are largely driven by benefits such as increased safety (through reduced human error), lower operating costs, and reduced emissions, all of which hold significant promise for reshaping shipping operations and logistics (gu2021autonomous; psaraftis2023logistics).
To bridge the gap between concept and full-scale deployment, researchers and industry leaders have developed or repurposed vessels of varying sizes as experimental platforms for autonomous navigation (schiaretti2017survey). The milliAmpere 1 (hinostroza2024_OMAE2025) and milliAmpere 2 (eide2025autonomous) ferries—both built at half-scale for urban waterways—have facilitated the development and validation of key autonomy functions such as sensor fusion, collision avoidance, and automated docking under real harbor conditions (martinsen2019autonomous). Yara Birkeland (Skredderberget2025) is a fully electric container ship providing a full-scale test platform for uncrewed cargo operations. Furthermore, the crewed research vessel R/V Gunnerus (marley2023four) is equipped with conventional ship systems and an extensive suite of sensors and data acquisition tools, and it provides a platform for advanced scientific studies. In the commercial scale, companies like Maritime Robotics and SeaRobotics each offer a fleet of autonomous surface vessels (ASVs) of multiple size for robotics applications and hydrographic surveys, while Reach Subsea, Ocean Infinity, and Fugro each develop their fleets of autonomous vessels specifically designed for offshore operations.
@@
In this work, we introduce a digital-physical testbed for Maritime Autonomous Surface Ships (MASS) that combines high-fidelity simulations, reduced-order models, and real-time sensor feedback within a unified framework. Our testbed facilitates a smooth transition from simulation to physical experimentation, supported by a digital twin environment in Unity for visualization and remote control. This setup aims to enhance the development and validation of autonomy algorithms by providing a cost-effective and flexible platform.
@@ 2.1 Overview
Cybership II skjetne2004modeling End of life
Cybership III nguyen2007design End of life
@@
Table 1: The C/S Fleet and research vessels at the time of writing.
To complement these physical assets, we have developed corresponding simulation environments for each vessel. These digital platforms support early-stage algorithm development and debugging, ranging from simple integration-test models to high-fidelity simulations with realistic hydrodynamic effects. Additionally, sensor-model simulations enable testing of algorithms that rely on complex sensor feedback. We also employ a digital twin environment in Unity, allowing us to visualize and analyze vessel behaviors, and gain better insights.
Taken together, these components form our design, verification, test, and validation pipeline. Algorithms are first iterated upon in simulation—where rapid testing reduces cost and risk—before being transferred to the physical model vessels in our laboratory basin. This progressive approach not only streamlines the overall testing process but also ensures that MASS algorithms are thoroughly verified and validated across increasing levels of realism and operational complexity prior to deployment in real maritime conditions.
@@ 3.1.1 High-fidelity simulation model
The high-fidelity simulation framework mcsimpy [^2] The mcsimpy open-source can be accessed at https://github.com/NTNU-MCS/mcsimpy, written in Python, replicates the dynamics of the C/S fleet by incorporating wave-excitation models and realistic hydrodynamic coefficients. Specifically, it uses JONSWAP (hasselmann1973measurements) and Pierson–Moskowitz (pierson1964proposed) wave spectra to generate diverse sea states, while hydrodynamic parameters derived from WAMIT (wamit7v5) and VERES (shipx). mcsimpy supports both 3 and 6 degree of freedom (DOF) models, making it suitable for various vessel configurations, including the R/V Gunnerus. Built on the studies by mo2023real; hygen2023deterministic and the Marine Systems Simulator perez2006overview, this framework provides a versatile platform for simulating control and navigation algorithms under realistic environmental conditions. This simulation environment is designed for control algorithm development, controller tuning, and serving as a fast, flexible sandbox for machine learning (ML) and reinforcement learning (RL) applications.
@@ 3.1.2 Reduced-order model simulation
We employ a lightweight simulation environment [^3] The simulator for the simplified hydrodynamic model is open-source and available at https://github.com/NTNU-MCS/shoeboxpy that approximates vessel dynamics using reduced-order models. This environment is primarily used for rapid testing of software interfaces and high-level control logic as it shares the same programming interface as the physical platform.
@@ 3.1.3 Simplified dynamics with sensor models
Stonefish (cieslak2019stonefish) is a high performance and real-time capable robotics simulator written in C++. Here, it is used for providing a sandbox for image processing, supports a wide range of sensor types (such as cameras, inertial measurement units, and sonars), and has simplified collision and wave dynamics. Thanks to its flexibility, Stonefish is ideal for evaluating algorithms like localization, object detection, and platform interaction that requires environmental variability and realistic sensor feedback. Stonefish bridges the gap between reduced-order simulation models and large-scale real-world tests by simulating important elements of the environment. This allows validation of control and perception algorithms prior to their deployment.
@@ 3.2 Digital Twin
The digital twin system is paired with a remote control center, which allows operators to control the physical vessels in the MC-Lab from a separate location (halvorsen2024combining). Command can be performed manually using joysticks or Metaquest 2 headset and hand controls [^4] Link to the video demonstration: https://www.youtube.com/watch?v=fqEdp8VXAqE, as well as using these interfaces to provide guidance commands like waypoints and speed references for automatic control. Vessel state information, such as position and orientation, is transmitted over the site-to-site network to provide remote monitoring capabilities. While some limitations exist due to hardware constraints—particularly in the responsiveness of the monitoring—the system has been tested in both simulation and physical experiments and has shown to function reliably under the available conditions.
@@ 3.3.2 Control and estimation modules
The C/S fleet shares a unified and modular control architecture that simplifies integration into the broader testbed infrastructure and provides a consistent foundation for algorithm development. Built-in control and estimation modules offer baseline performance metrics, serving as reference implementations for testing and benchmarking new strategies. Depending on the experimental need, these modules can be disabled or replaced with custom implementations, allowing researchers to test their own algorithms in a consistent environment. To this end, we choose to keep the default modules as simple as possible, while still providing the necessary functionality for basic control and estimation tasks. The overall control architecture is illustrated in Fig. 3. Using this architecture, we can easily test and validate various control algorithms in a consistent manner.
@@
A proportional-integral reference feed-forward (PI-RFF) velocity controller is made available to track desired velocity. For the body-frame velocity $\nu:=[u,v,r]^{\top}\in\mathbb{R}^{3}$, the 3-DOF surge–sway–yaw plant without Coriolis or centrifugal terms is
@@
with $\mathbf{M}=\mathbf{M}^{\top}>0$ and $\mathbf{D}\succ 0$. Let the reference velocity be $\nu_{d}(t)$, define the tracking error $\tilde{\nu}:=\nu-\nu_{d}$, the integral state $\dot{\xi}=\tilde{\nu}$, and choose $a$ as the virtual control input,
@@
where $K_{p}=\operatorname{diag}(k_{p1},k_{p2},k_{p3})\succ 0$ and $K_{i}=\operatorname{diag}(k_{i1},k_{i2},k_{i3})\succ 0$. The commanded forces/torques become
@@
which combines feed-forward $\mathbf{M}\dot{\nu}_{d}$, proportional-integral feedback, and compensates for damping using $\mathbf{D}\nu_{d}$. Consider the Lyapunov function,
$V(\tilde{\nu},\xi)=\tfrac{1}{2}\tilde{\nu}^{\top}\mathbf{M}\tilde{\nu}+\tfrac{1}{2}\xi^{\top}K_{i}\xi,$ (7)
@@
Because $\mathbf{D}+\mathbf{M}K_{p}\succ 0$, the Lyapunov function $V$ is positive-definite and radially unbounded, while $\dot{V}$ is negative semi-definite. Hence, the control law (6) guarantees stability of the closed-loop system.
@@
We employ a mass-damper-spring based reference model for stationkeeping controller. The reference filter defines a state vector $\mathbf{x}(t)=[\eta^{\top}_{d}(t),\dot{\eta}^{\top}_{d}(t),\ddot{\eta}^{\top}_{d}(t)]^{\top}$, where $\eta_{d}\in\mathbb{R}^{3}$ is the desired pose in the world frame. The filter evolves according to the continuous-time state-space model
@@
with $\mathbf{\Omega}=\mathrm{diag}(\omega_{1},\omega_{2},\omega_{3})$ controlling the filter bandwidth, and $\mathbf{\Delta}=\mathrm{diag}(\delta_{1},\delta_{2},\delta_{3})$ controlling the filter damping (fossen2021handbook). Other types of reference models may also be implemented, see (fernandes2012reference). A forward-Euler discretization step is sufficient to update the filter state, providing smooth desired pose, velocity, and acceleration outputs for the DP controller. The default included controller is simply a PD control law to generate the desired thrust loads for the vessel to closely follow the reference trajectory using
@@
The four corner test (skjetne2017amos) conducted using the pose controller and the results in Fig. 4 shows the performance differences between simulation and physical implementation. The simulation achieved a position RMSE of 0.062 m, yaw RMSE of $0.439^{\circ}$, and velocity RMSE of 0.031 m/s, whereas the physical vessel showed 0.180 m, $4.295^{\circ}$, and 0.036 m/s, respectively.
@@
A multiplexer is employed to dynamically select between different control modes. This enables operators to choose the most appropriate control strategy for a given task on the fly.
@@ 4 Physical infrastructure
The MC-Lab provides a versatile, well-instrumented environment for testing model-scale marine vessels in realistic wave conditions. Its primary facility is a wave basin measuring $32m$ from the flap-type wave maker to the beach and 6.4m wide, with a maximum depth of $1.5m$. The wave maker flap spans the full basin width, and is capable of generating waves up to $0.25m$ in amplitude at controlled frequencies, and a wave-damping beach is located at the far end that minimizes reflections for cleaner experiments. This allows users to produce both regular and irregular wave spectra, such as JONSWAP, in either manual or automated modes. Conductive and ultrasonic probes can be places throughout the basin to monitor wave characteristics. A Motion Capture system offers real-time object tracking (to within $1cm$ in position and $0.5^{\circ}$ in orientation). Additional technical specifications and usage details can be found in the web-pages of MC-lab. By 2030, the MC-Lab will be succeeded by a new model basin for academic research, called the Hydro-Cybernetics Laboratory in the House of Archimedes as the part of the Norwegian Ocean Technology Centre (NTNU_Norwegian_Ocean_Technology_Centre). This more modern facility will be a larger test basin, approximately $65m$ in length, $8m$ in width, and $3m$ in depth, further expanding research capabilities in hydrodynamics, marine cybernetics, and ocean technology.
@@ 4.1 Hardware
All vessels in the C/S fleet share a standardized hardware architecture designed for modularity, scalability, and ease of integration. The primary differences between the vessels is the number and types of actuators and sensors, which are abstracted through a common software stack. This allows all vessels to operate with minimal configuration differences, streamlining development and deployment across the fleet. As a result, new vessels can be added to the testbed environment with minimal re-engineering, using the existing architecture as a template for both hardware and software. This consistency reduces development time and increases maintainability, reliability, and scalability.
At the core of each vessel is a Raspberry Pi single-board computer, which serves as the main onboard processor. This choice offers a compact, capable, and well documented platform for running control algorithms, handling sensor input, and managing actuator commands in real time. Communication between onboard components is kept streamlined and maintainable by using standard hardware protocols, ensuring compatibility with a wide range of off-the-shelf sensors and devices, see Fig. 5.

### v2

@@ Abstract
The algorithms developed for Maritime Autonomous Surface Ships (MASS) are often challenging to test on actual vessels due to high operational costs and safety considerations. Simulations offer a cost-effective alternative and eliminate risks, but they may not accurately represent real-world dynamics for the given tasks. Utilizing small-scale model ships and robotic vessels in conjunction with a laboratory basin provides an accessible testing environment for the early stages of validation processes. However, designing and developing a model vessel for a single test can be costly and cumbersome, and researchers often lack access to such infrastructure. To address these challenges and enable streamlined testing, we have developed an in-house testbed that facilitates the development, testing, verification, and validation of MASS algorithms in a digital-physical laboratory. This infrastructure includes a set of small-scale model vessels, a simulation environment for each vessel, a comprehensive testbed environment, and a digital twin in Unity. With this, we aim to establish a full design and verification pipeline that starts with high-fidelity simulation models of each model vessel, to the model-scale testing in the laboratory basin, allowing possibilities for moving towards semi-fullscale validation with R/V milliAmpere1 and full-scale validation with R/V Gunnerus. In this work, we present our progress on the development of this testbed environment and its components, demonstrating its effectiveness in enabling ship guidance, navigation, and control (GNC), including autonomy.
@@ 1 Introduction
Autonomous vessels have the potential to significantly influence the maritime sector by enhancing safety, operational efficiency, and economic viability (gu2021autonomous; munim2019autonomous). Recent advances in sensor technology, artificial intelligence, and marine cybernetics have spurred the development of prototype autonomous vessels of various sizes (schiaretti2017survey). Additionally, large-scale initiatives have emerged to demonstrate the real-world feasibility of uncrewed ships under diverse conditions (reddy2019zero). These efforts are largely driven by benefits such as increased safety (through reduced human error), lower operating costs, and reduced emissions, all of which hold significant promise for reshaping shipping operations and logistics (gu2021autonomous; psaraftis2023logistics).
To bridge the gap between concept and full-scale deployment, researchers and industry leaders have developed or repurposed vessels of varying sizes as experimental platforms for autonomous navigation (schiaretti2017survey). The milliAmpere1 (hinostroza2024_OMAE2025) and milliAmpere2 (eide2025autonomous) ferries—both built at half-scale for urban waterways—have facilitated the development and validation of key autonomy functions such as sensor fusion, collision avoidance, and automated docking under real harbor conditions (martinsen2019autonomous). Yara Birkeland (Skredderberget2025) is a fully electric container ship providing a full-scale test platform for uncrewed cargo operations. Furthermore, the crewed research vessel R/V Gunnerus (marley2023four) is equipped with conventional ship systems and an extensive suite of sensors and data acquisition tools, and it provides a platform for advanced scientific studies. In the commercial scale, companies like Maritime Robotics and SeaRobotics each offer a fleet of autonomous surface vessels (ASVs) of multiple sizes for robotics applications and hydrographic surveys, while Reach Subsea, Ocean Infinity, and Fugro each develop their fleets of autonomous vessels specifically designed for offshore operations.
@@
In this work, we introduce a digital-physical testbed for Maritime Autonomous Surface Ships (MASS) that combines high-fidelity simulations, reduced-order models, and real-time sensor feedback within a unified framework. Our testbed facilitates a smooth transition from simulation to physical experimentation, supported by a digital twin environment in Unity for visualization and remote control. This setup aims to enhance the development and validation of autonomy algorithms by providing a cost-effective and flexible platform. The primary contribution is an open, modular software-hardware stack—featuring curated datasets and a validated hydrodynamic model library—that transforms the Marine Cybernetics Laboratory basin into a benchmark platform for ship autonomy research.
@@ 2.1 Overview
Cybership II skjetne2005maneuvering End of life
Cybership III smogeli2006control End of life
@@
Table 1: The C/S Fleet and research vessels at the time of writing with relevant references.
To complement these physical assets, we have developed corresponding simulation environments for each vessel. These digital platforms support early-stage algorithm development and debugging, ranging from simple integration-test models to high-fidelity simulations with realistic hydrodynamic effects. Additionally, sensor-model simulations enable testing of algorithms that rely on complex sensor feedback. We also employ a digital twin environment in Unity, allowing us to visualize and analyze vessel behaviors and gain better insights.
Taken together, these components form our design, verification, test, and validation pipeline. Algorithms are first iterated upon in simulation, where rapid testing reduces cost and risk, before being transferred to the physical model vessels in our laboratory basin. This progressive approach not only streamlines the overall testing process but also ensures that MASS algorithms are thoroughly verified and validated across increasing levels of realism and operational complexity prior to deployment in real maritime conditions.
@@ 3.1.1 High-fidelity simulation model
The high-fidelity simulation framework mcsimpy [^2] The mcsimpy open-source can be accessed at https://github.com/NTNU-MCS/mcsimpy, written in Python, replicates the dynamics of the C/S fleet by incorporating wave-excitation models and realistic hydrodynamic coefficients. Specifically, it uses JONSWAP (hasselmann1973measurements) and Pierson–Moskowitz (pierson1964proposed) wave spectra to generate diverse sea states, while hydrodynamic parameters are derived from WAMIT (wamit7v5) and VERES (shipx). mcsimpy supports both 3 and 6 degrees of freedom (DOF) models, making it suitable for various vessel configurations, including the R/V Gunnerus. Built on the studies by mo2023real,hygen2023deterministic as well as the Marine Systems Simulator (perez2006overview), this framework provides a versatile platform for simulating control and navigation algorithms under realistic environmental conditions. The intended use of the simulation environment is for control algorithm development, controller tuning, and serving as a fast, flexible sandbox for machine learning (ML) and reinforcement learning (RL) applications.
@@ 3.1.2 Reduced-order model simulation
We employ a lightweight simulation environment [^3] The simulator for the simplified hydrodynamic model is open-source and available at https://github.com/NTNU-MCS/shoeboxpy that approximates vessel dynamics using reduced-order models. This environment is primarily used for rapid testing of software interfaces and high-level control logic, as it shares the same programming interface as the physical platform.
@@ 3.1.3 Simplified dynamics with sensor models
Stonefish (cieslak2019stonefish) is a high performance and real-time capable robotics simulator written in C++. Here, it is used for providing a sandbox for image processing, supports a wide range of sensor types (such as cameras, inertial measurement units, and sonars), and has simplified collision and wave dynamics. Thanks to its flexibility, Stonefish is ideal for evaluating algorithms like localization, object detection, and platform interaction that requires environmental variability and realistic sensor feedback. Stonefish bridges the gap between reduced-order simulation models and large-scale real-world tests by simulating important elements of the environment, and allows validation of control and perception algorithms.
@@ 3.2 Digital Twin
The digital twin system is paired with a remote control center, which allows operators to control the physical vessels in the MC-Lab from a separate location (halvorsen2024combining). Command can be performed manually using joysticks or Metaquest 2 headset and hand controls [^4] Link to the video demonstration: https://www.youtube.com/watch?v=fqEdp8VXAqE, as well as using these interfaces to provide guidance commands like waypoints and speed references for automatic control. Vessel state information, such as position and orientation, is transmitted over the site-to-site network to provide remote monitoring capabilities. While some limitations exist due to hardware constraints—particularly in the responsiveness of the monitoring—the system has been tested in both simulation and physical experiments and has been shown to function reliably under the available conditions.
@@ 3.3.2 Control and estimation modules
The C/S fleet shares a unified and modular control architecture that simplifies integration into the broader testbed infrastructure and provides a consistent foundation for algorithm development. Built-in control and estimation modules offer baseline performance metrics, serving as reference implementations for testing and benchmarking new strategies. Depending on the experimental need, these modules can be disabled or replaced with custom implementations, allowing researchers to test their own algorithms in a consistent environment. To this end, we choose to keep the default modules as simple as possible, while still providing the necessary functionality for basic control and estimation tasks. The overall control architecture is illustrated in Fig. 3. Using this architecture, we can easily test and validate various control algorithms consistently.
@@
A proportional-integral with reference feedforward (PI-RFF) velocity controller is made available to track a desired velocity. For the body-frame velocity $\nu:=[u,v,r]^{\top}\in\mathbb{R}^{3}$, the 3-DOF surge–sway–yaw plant without Coriolis or centrifugal terms is
@@
where $\mathbf{M}=\mathbf{M}^{\top}>0$ and $\mathbf{D}>0$. Let the reference velocity be $\nu_{d}(t)$, define the tracking error $\tilde{\nu}:=\nu-\nu_{d}$, the integral state $\dot{\xi}=\tilde{\nu}$, and choose $a$ as the virtual control input,
@@
where $K_{p}=\operatorname{diag}(k_{p1},k_{p2},k_{p3})>0$ and $K_{i}=\operatorname{diag}(k_{i1},k_{i2},k_{i3})>0$. The commanded loads (forces/torques) become
@@
which combines acceleration reference feedforward $\mathbf{M}\dot{\nu}_{d}$, proportional-integral feedback, and feedforward compansation of damping using $\mathbf{D}\nu_{d}$. Consider the Lyapunov function,
$V(\tilde{\nu},\xi)=\tfrac{1}{2}\tilde{\nu}^{\top}\mathbf{M}\tilde{\nu}+\tfrac{1}{2}\xi^{\top}K_{i}\mathbf{M}\xi,$ (7)
@@
Because $\mathbf{D}+\mathbf{M}K_{p}>0$, the Lyapunov function $V$ is positive-definite and radially unbounded, while $\dot{V}$ is negative semi-definite. A straightforward application of the Krasovskii-LaSalle invariance principle will show that the equilibrium $(\tilde{\nu},\xi)=(0,0)$ is uniformly globally asymptotically stable.
@@
We employ a mass-damper-spring based reference model for the stationkeeping controller. The reference filter defines a state vector $\mathbf{x}(t)=[\eta^{\top}_{d}(t),\dot{\eta}^{\top}_{d}(t),\ddot{\eta}^{\top}_{d}(t)]^{\top}$, where $\eta_{d}\in\mathbb{R}^{3}$ is the desired pose in the world frame. The filter evolves according to the continuous-time state-space model
@@
with $\mathbf{\Omega}=\mathrm{diag}(\omega_{1},\omega_{2},\omega_{3})$ controlling the filter bandwidth, and $\mathbf{\Delta}=\mathrm{diag}(\delta_{1},\delta_{2},\delta_{3})$ controlling the filter damping (fossen2021handbook). Other types of reference models may also be implemented; see (fernandes2012reference). A forward-Euler discretization step is sufficient to update the filter state, providing smooth desired pose, velocity, and acceleration outputs for the DP controller. The default included controller is simply a PD control law to generate the desired thrust loads for the vessel to closely follow the reference trajectory using
@@
The 4-corner test (skjetne2017amos) was conducted using the pose controller, and the results in Fig. 4 show the performance differences between simulation and physical implementation. The simulation achieved a position RMSE of 0.062m, yaw RMSE of $0.439^{\circ}$, and velocity RMSE of 0.031m/s, whereas the physical vessel showed 0.180m, $4.295^{\circ}$, and 0.036m/s, respectively.
@@
A multiplexer is used to choose between different control modes. This enables operators to choose the most appropriate control strategy for a given task on the fly.
@@ 4 Physical infrastructure
The MC-Lab provides a versatile and well-instrumented environment for testing model-scale marine vessels in realistic wave conditions. Its primary facility is a wave basin measuring $32m$ from the flap-type wave maker to the beach and 6.4m wide, with a maximum depth of $1.5m$. The wave maker flap spans the full basin width, and is capable of generating waves up to $0.25m$ in amplitude at controlled frequencies, and a wave-damping beach is located at the far end that minimizes reflections for cleaner experiments. This allows users to produce both regular and irregular wave spectra, such as JONSWAP, in either manual or automated modes. Conductive and ultrasonic probes can be placed throughout the basin to monitor wave characteristics. A Motion Capture system offers real-time object tracking (to within $1cm$ in position and $0.5^{\circ}$ in orientation). Additional technical specifications and usage details can be found on the web pages of MC-lab. Within 2030, the MC-Lab will be succeeded by a new model basin for academic research, called the Hydro-Cybernetics Laboratory in the House of Archimedes as part of the new Norwegian Ocean Technology Centre (NTNU_Norwegian_Ocean_Technology_Centre). This more modern facility will be a larger test basin, approximately $65m$ in length, $8m$ in width, and $3m$ in depth, further expanding research capabilities in hydrodynamics, marine cybernetics, and ocean technology.
@@ 4.1 Hardware
All vessels in the C/S fleet share a standardized hardware architecture designed for modularity, scalability, and ease of integration. The primary differences between the vessels are the number and types of actuators and sensors, which are abstracted through a common software stack. This allows all vessels to operate with minimal configuration differences, streamlining development and deployment across the fleet. As a result, new vessels can be added to the testbed environment with minimal re-engineering, using the existing architecture as a template for both hardware and software. This consistency reduces development time and increases maintainability, reliability, and scalability.
At the core of each vessel is a Raspberry Pi single-board computer, which serves as the main onboard processor. This choice offers a compact, capable, and well documented platform for running control algorithms, handling sensor input, and managing actuator commands in real time. Communication between onboard components is kept streamlined and maintainable by using standard hardware protocols, ensuring compatibility with a wide range of off-the-shelf sensors and devices; see Fig. 5.